# Time-Series at the Edge: Tiny Separable CNNs for Wearable Gait Detection and Optimal Sensor Placement

### v1

@@ Conclusion
This work evaluated 1D convolutional neural network architectures and a threshold baseline for gait-event detection in people with Parkinson’s disease using accelerometer data from different body locations. It was shown how chest data clearly outperformed other sensor positions, which included arms and legs. The goal is to detect gait windows on-sensor to replace raw data streams to improve bandwidth, storage, and energy. In leave-one-subject-out validation, all CNNs outperformed the threshold baseline, which in general shows poor generalization. A detailed deployment analysis for embedded systems demonstrated that all models run within memory and latency constraints of representative STM32 microcontrollers. Model 2, which adds residual connections on top of the separable-convolution design of Model 1 and the baseline, achieves the best accuracy–efficiency trade-off, with sub-millisecond inference on H7 and 5–7,ms on L4/F4. Model 1, which employs a simpler separable-convolution architecture, is a valid alternative with slightly lower F1. Fixed-threshold detectors, while extremely cheap, are sensitive to inter-subject variation and under-perform on certain positions relative to CNN models.

### v2

@@ Conclusion
This work evaluated 1D convolutional neural network architectures and a threshold baseline for gait-event detection in people with Parkinson’s disease using accelerometer data from different body locations. It was shown how chest data clearly outperformed other sensor positions, which included arms and legs. The goal is to detect gait windows on-sensor to replace raw data streams to improve bandwidth, storage, and energy. In leave-one-subject-out validation, all CNNs outperformed the threshold baseline, which in general shows poor generalization. A detailed deployment analysis for embedded systems demonstrated that all models run within memory and latency constraints of representative STM32 microcontrollers. Model 2, which adds residual connections on top of the separable-convolution design of Model 1 and the baseline, achieves the best accuracy–efficiency trade-off, with sub-millisecond inference on H7 and 5–7 ms on L4/F4. Model 1, which employs a simpler separable-convolution architecture, is a valid alternative with slightly lower F1. Fixed-threshold detectors, while extremely cheap, are sensitive to inter-subject variation and under-perform on certain positions relative to CNN models.